# Looped Transformers are Better at Learning Learning Algorithms

### v2

@@ 2 Related Works
In addition to approaching this problem from an empirical perspective, several studies have examined it from a theoretical standpoint, focusing on the construction of transformers and the linear attention network (excluding softmax counterparts). They connected the learnability of the transformer for the linear regression task with the implicit gradient descent updates on query prompt predictions (von Oswald et al., 2022; Ahn et al., 2023; Zhang et al., 2023a; Mahankali et al., 2023; Ding et al., 2023). von Oswald et al. (2022) empirically demonstrated that when minimizing the $\ell_{2}$ distance between predictions and true labels, one step of gradient descent is the optimal solution for a transformer to learn when solving the linear regression problem, while Ahn et al. (2023); Zhang et al. (2023a) and Mahankali et al. (2023) provided theoretical explanations for it. Moreover, another line of work (Muller et al., 2021; Xie et al., 2021; Akyürek et al., 2022; Bai et al., 2023) provided a theoretical understanding of the transformer’s in-context learnability through the lens of the Bayesian estimator. Li et al. (2023) studied the in-context learning from the viewpoint of generalization error.
@@ 4 Training Algorithm for Looped Transformer
For comparison with the unlooped transformer, we use a transformer identical to the one described in Garg et al. (2022), except for the number of layers. Specifically, we employ a GPT-2 model with an embedding dimension of $D=256$ and $h=8$ attention heads. The standard (unlooped) transformer has $L=12$ layers, and the looped transformer has $L=1$ layer. In terms of the number of parameters, the unlooped transformer comprises 9.489.48M parameters [^2] The discrepancy in the number of parameters reported here and in Garg et al. (2022) (22.422.4M) arises because they also include the word-to-embedding parameters, which are not utilized in the model. Hence, we exclude these parameters from the overall count., and the looped transformer uses 0.790.79M. We follow the same training strategy: train with Adam optimizer, learning rate $0.0001$, no weight decay or other explicit regularization (such as gradient clip, or data augmentation).
@@ 5.1 Experimental Setup
Consequently, we avoid referring to this approach as “curriculum learning,” as that term typically refers to starting with a less difficult task and gradually moving on to more challenging ones. Rather, this strategy can be seen as a method to kick-start the training of the transformer model in a loop structure. In general, the decision to use or not to use the scheduling does not significantly impact the outcome. However, in some instances, we observe superior results achieved by training with a schedule for $b$. More details can be found in Appendix D.
@@ 5.2 Looped Transformer can in-context Learn Linear Regression
In the task of (b) noisy linear regression, the looped transformer displays a reduced peak in the double descent curve, similar to the effects of applying minor regularization to the base model for resolving noisy linear regression. Bhattamishra et al. (2022) also reported the simplicity bias of the random transformer towards low sensitivity in the boolean function base. However, this inductive bias towards simplicity can negatively impact performance when there is a scaling shift in the input distribution. As depicted in Fig. 5 (c), during the inference process, if we sample $\bm{x}$} such that ${\bm{x}}=3{\bm{z}}$, and ${\bm{z}}\sim\mathcal{N}(0,I)$, the looped transformer underperforms compared to the unlooped transformer. More extrapolation results can be found in Appendix C.
@@ 5.3 Impact of Model Architecture Variations on Looped Transformer Performance
To delve deeper into the learning dynamics of each layer in both the transformer and the looped transformer, we employ the model probing technique suggested by Akyürek et al. (2022); Alain & Bengio (2016). Reusing some notation, we represent the output of the standard transformer’s $t$-th layer or the looped transformer’s $t$ loop iterations as $Y_{t}$. An MLP model is trained on $Y_{t}$ to learn target probes, specifically ${\bm{X}}^{T}{\bm{y}}$ for the gradient component and the ${\bm{w}}_{ols}$ for the optimal least squares solution. The mean squared error for model probing is illustrated in Fig. 7. While standard transformers initially capture representations relevant to ${\bm{X}}^{T}{\bm{y}}$ and ${\bm{w}}_{ols}$, and subsequently shift focus to other statistics, the looped transformer consistently refines its representations related to the target probes across its loop iterations. Details of the model probing are presented in Appendix E.
@@
Fig. 6 (right) illustrates the squared error of the looped transformer for various combinations of $L$, $D$, and $h$. With $L=1$ and $h=8$ held constant, increasing $D$ leads to improved convergence, saturating at $D=256$. When holding $D=256$ and $h=8$ constant and varying $L$, the looped transformer shows similar convergence performance. The primary differences are in the speed of convergence. Adjusting only $h$ with fixed $L$ and $D$, the looped transformer’s error decreases as $h$ rises from 2 to 8. However, the error increases when $h$ progresses from 8 to 32. Here, we follow the standard implementation of GPT-2333https://github.com/karpathy/nanoGPT, where the per-head embedding dimension is defined as $D\over h$. This definition implies a trade-off in the transformer between the number of heads and the per-head embedding dimension as $h$ varies.
@@ 5.4 Higher Complexity Function Classes
In this section, we shift our focus to function classes of higher complexity: a) the sparse linear model with $d=20$, $k=100$, and non-sparse entry $s=3$; b) the decision tree with depth=4=4 and input dimension $d=20$; c) the 2-layer ReLU neural network with $100$ hidden neurons and input dimension $d=20$. For these functions, parameters are sampled from $\mathcal{N}(0,I_{d})$, and in-context samples ${\bm{x}}_{i}\sim\mathcal{N}(0,I_{d})$. This setup follows the methodology described in Garg et al. (2022), and is further detailed in Appendix A. We also conduct experiments on OpenML datasets (Vanschoren et al., 2013), which better represents practical scenarios. Details of experiment setup and results for OpenML datasets are available in Appendix F.
@@ 6 Discussion and Future Work
The expressive power of recurrent models has been widely investigated in the literature: Theorem 3 in Bai et al. (2019) demonstrated that any traditional $L$-layer neural network can be represented by a weight-tied, input-injected network. Further, Giannou et al. (2023) explored the potential of an encoder looped transformer to function as a universal computer. Our work, instead, focuses on the empirical validation of this looped architecture, and presents a practical training method to effectively train a looped decoder transformer for in-context learning. Establishing the theorem regarding the expressiveness of the looped decoder transformer presents an interesting topic for future research.
@@ Appendix B Choice of Loop Iterations and its Effect on In-context Performance
Appendix C Extrapolation Behavior of Looped Transformer and Unlooped Transformer
@@
Appendix D Effects of Scheduling
@@
Appendix E Details of Model Probing
@@
Appendix F OpenML Experiment

### v3

@@ 2 Related Works
In addition to approaching this problem from an empirical perspective, several studies have examined it from a theoretical standpoint, focusing on the construction of transformers and the linear attention network (excluding softmax counterparts). They connected the learnability of the transformer for the linear regression task with the implicit gradient descent updates on query prompt predictions (von Oswald et al., 2022; Ahn et al., 2023; Zhang et al., 2023a; Mahankali et al., 2023; Ding et al., 2023). von Oswald et al. (2022) empirically demonstrated that when minimizing the $\ell_{2}$ distance between predictions and true labels, one step of gradient descent is the optimal solution for a one-layer linear self-attention (LSA) transformer to learn when solving the linear regression problem, while Ahn et al. (2023); Zhang et al. (2023a; 2024) and Mahankali et al. (2023) provided theoretical explanations for it. Beyond 1-layer LSA, Fu et al. (2023); Giannou et al. (2024) have demonstrated that multi-layer transformers can perform higher-order optimization. Moreover, another line of work (Muller et al., 2021; Xie et al., 2021; Akyürek et al., 2022; Bai et al., 2023) provided a theoretical understanding of the transformer’s in-context learnability through the lens of the Bayesian estimator. Li et al. (2023) studied the in-context learning from the viewpoint of generalization error. Lin & Lee (2024) investigated the dual operating modes of in-context learning.
@@ 4 Training Algorithm for Looped Transformer
For comparison with the unlooped transformer, we use a transformer identical to the one described in Garg et al. (2022), except for the number of layers. Specifically, we employ a GPT-2 model with an embedding dimension of $D=256$ and $h=8$ attention heads. The standard (unlooped) transformer has $L=12$ layers, and the looped transformer has $L=1$ layer. In terms of the number of parameters, the unlooped transformer comprises 9.489.48M parameters, and the looped transformer uses 0.790.79M. We follow the same training strategy: train with Adam optimizer, learning rate $0.0001$, no weight decay or other explicit regularization (such as gradient clip, or data augmentation).
@@ 5.1 Experimental Setup
Consequently, we avoid referring to this approach as “curriculum learning,” as that term typically refers to starting with a less difficult task and gradually moving on to more challenging ones. Rather, this strategy can be seen as a method to kick-start the training of the transformer model in a loop structure. In general, the decision to use or not to use the scheduling does not significantly impact the outcome. However, in some instances, we observe superior results achieved by training with a schedule for $b$. More details can be found in Appendix E.
@@ 5.2 Looped Transformer can in-context Learn Linear Regression
Looped Transformer Learns Efficiently with Fewer Distinct In-Distribution Prompts.
We further evaluate the sample complexity of transformer training. Specifically, we ask how many distinct prompts need to be seen during training for the transformer or the looped transformer to learn the function. To do so, instead of generating the linear function on the fly in each iteration, we generate the training dataset before training. Thus, during training, the transformer may encounter the same prompt multiple times.
To isolate the impact of curriculum learning, we disable this strategy during and focus on the linear regression task with problem dimension $d=10$, in-context samples $k=20$. The results are presented in Fig. 5. Due to the fewer parameters in the looped transformer, it is able to learn the function with fewer distinct prompts/functions compared to the standard transformer. More results regarding different problem dimensions are presented in Appendix C.
Figure 5: Performance of transformer on linear functions with $d=10$ and $k=21$, when trained with different numbers of distinct prompts/functions.
@@
In the task of (b) noisy linear regression, the looped transformer displays a reduced peak in the double descent curve, similar to the effects of applying minor regularization to the base model for resolving noisy linear regression. Bhattamishra et al. (2022) also reported the simplicity bias of the random transformer towards low sensitivity in the boolean function base. However, this inductive bias towards simplicity can negatively impact performance when there is a scaling shift in the input distribution. As depicted in Fig. 6 (c), during the inference process, if we sample $\bm{x}$} such that ${\bm{x}}=3{\bm{z}}$, and ${\bm{z}}\sim\mathcal{N}(0,I)$, the looped transformer underperforms compared to the unlooped transformer. More extrapolation results can be found in Appendix D.
@@ 5.3 Impact of Model Architecture Variations on Looped Transformer Performance
To delve deeper into the learning dynamics of each layer in both the transformer and the looped transformer, we employ the model probing technique suggested by Akyürek et al. (2022); Alain & Bengio (2016). Reusing some notation, we represent the output of the standard transformer’s $t$-th layer or the looped transformer’s $t$ loop iterations as $Y_{t}$. An MLP model is trained on $Y_{t}$ to learn target probes, specifically ${\bm{X}}^{T}{\bm{y}}$ for the gradient component and the ${\bm{w}}_{ols}$ for the optimal least squares solution. The mean squared error for model probing is illustrated in Fig. 8. While standard transformers initially capture representations relevant to ${\bm{X}}^{T}{\bm{y}}$ and ${\bm{w}}_{ols}$, and subsequently shift focus to other statistics, the looped transformer consistently refines its representations related to the target probes across its loop iterations. Details of the model probing are presented in Appendix F.
@@
Fig. 7 (right) illustrates the squared error of the looped transformer for various combinations of $L$, $D$, and $h$. With $L=1$ and $h=8$ held constant, increasing $D$ leads to improved convergence, saturating at $D=256$. When holding $D=256$ and $h=8$ constant and varying $L$, the looped transformer shows similar convergence performance. The primary differences are in the speed of convergence. Adjusting only $h$ with fixed $L$ and $D$, the looped transformer’s error decreases as $h$ rises from 2 to 8. However, the error increases when $h$ progresses from 8 to 32. Here, we follow the standard implementation of GPT-2222https://github.com/karpathy/nanoGPT, where the per-head embedding dimension is defined as $D\over h$. This definition implies a trade-off in the transformer between the number of heads and the per-head embedding dimension as $h$ varies.
@@ 5.4 Higher Complexity Function Classes
In this section, we shift our focus to function classes of higher complexity: a) the sparse linear model with $d=20$, $k=100$, and non-sparse entry $s=3$; b) the decision tree with depth=4=4 and input dimension $d=20$; c) the 2-layer ReLU neural network with $100$ hidden neurons and input dimension $d=20$. For these functions, parameters are sampled from $\mathcal{N}(0,I_{d})$, and in-context samples ${\bm{x}}_{i}\sim\mathcal{N}(0,I_{d})$. This setup follows the methodology described in Garg et al. (2022), and is further detailed in Appendix A. We also conduct experiments on OpenML datasets, which better represents practical scenarios. Details of experiment setup and results for OpenML (Vanschoren et al., 2013) datasets are available in Appendix G.
@@ 6 Discussion and Future Work
The expressive power of recurrent models has been widely investigated in the literature: Theorem 3 in Bai et al. (2019) demonstrated that any traditional $L$-layer neural network can be represented by a weight-tied, input-injected network. Further, Giannou et al. (2023) explored the potential of an encoder looped transformer to function as a universal computer. Our work concentrates on empirically validating the looped architecture, offering a practical training method for a looped decoder transformer in in-context learning. Investigating the expressiveness of this transformer presents a promising future research direction.
@@ Appendix B Choice of Loop Iterations and its Effect on In-context Performance
Appendix C Effect of Number of Distinct Prompts / Functions Seen During Training
In Section 5, we investigated the impact of the number of distinct prompts during transformer training for a problem dimension $d=10$ and in-context samples $k=20$. In this section, we extend the results to include (1) $d=5$, $k=10$, and (2) $d=20$, $k=40$. It is important to note that all experiments were conducted without the use of curriculum learning for $d$ and $k$. For the looped transformer, we set $b=30$, and $T=15$, and we don’t apply scheduling in the loop iterations as well. For $d=5$, we investigated learning rate in $\{0.0001\}$, for $d=10$ with learning rate $\{0.00001,0.0001\}$, and for $d=20$, we explored learning rate $\{0.00001,0.0001,0.001\}$, selecting the result with the best performance. The result is presented in Fig. 14. Across different problem dimensions, the looped transformer demonstrated better sample efficiency compared to the vanilla transformer.
(a) Problem dimension 5
(b) Problem dimension 10
(c) Problem dimension 20
Figure 14: Performance of the standard transformer and the looped transformer on linear regression problem with problem dimension $d=5,10,20$, when trained with a different number of distinct prompts/functions.
Appendix D Extrapolation Behavior of Looped Transformer and Unlooped Transformer
@@
Appendix E Effects of Scheduling
@@
Appendix F Details of Model Probing
@@
Appendix G OpenML Experiment